# Nav-SCOPE: Swarm Robot Cooperative Perception and Coordinated Navigation

### v2

@@ III-B3 Fusion optimization
where $P$ is the probability of exploration to the left, and $Z$ is a normalization factor. The calculation is scalar, since the robots’ observations are mapped before into the probability domain. The decision is based on the principle of Maximum A Posteriori (MAP) to address the subtle slope of $\bm{\sigma}[\cdot]$. The fusion process expands the field of view of individual robots, thus significantly reducing the uncertainty of environmental observations. Consequently, redundant paths are eliminated by convergence or divergence effects in terms of robot paths.
IV Model Deployment
This section explains the planning process, the determination of parameters and variables.
IV-A Planning Process
The planning process based on Nav-SCOPE is a minimal reactive solution, which has proved to be versatile in a wide range of robots [8, 3]. Robots will go through a series of states when planning towards the target. A finite state machine (FSM) is illustrated in Fig. 3. When a robot $R_{a}$ observes no obstacles within its navigation scope, it will advance straight towards the target. In order to avoid obstacles for safe and automatic navigation, we introduce a repulsion force $\vec{F}_{aB}$. Given the distance vector $\vec{r}_{aB}$ to an obstacle $B$, $\vec{F}_{aB}$ is defined as a negative gradient of a monotonically increasing potential:
$V_{aB}\left(\left\|\vec{r}_{aB}\right\|\right)$ $=-\frac{V_{0B}}{\left\|\vec{r}_{aB}\right\|},$ (12)
$\vec{F}_{aB}\left(\left\|\vec{r}_{aB}\right\|\right)$ $=-\nabla_{\vec{e}_{aB}}V_{aB}=-\frac{V_{0B}}{\left\|\vec{r}_{aB}\right\|^{2}}\vec{e}_{aB},$
where $\vec{e}_{aB}$ is a unit vector towards the border of obstacle $B$, and $V_{0B}$ is a constant. Consider a case where the obstacle $B$ appears on the following path of $R_{a}$ within the planning distance $l_{th}$, then the robot $R_{a}$ should stop to maintain a safe distance. At this point, $\|\vec{r}_{aB}\|=l_{th}$, $\vec{F}_{aB}$ needs to balance the propulsion $\vec{F}_{a}$ in the direction of $\vec{e}_{a}$ to avoid danger:
$V_{0B}=\frac{l_{th}^{2}}{\cos{(\frac{\alpha}{2})}}\|\vec{F}_{a}\|.$ (13)
The reactive strategy of Nav-SCOPE diverges from PF methods [8, 1] in the following aspects:
  • Only the risk point within the navigation scope will be the source of the repulsion fields.
  • Obstacles in the peripheral view do not contribute to $\vec{F}_{aB}$.
  • Repulsive effects are considered independently.
The former two aspects reduce computation, and the third one avoids saddle points commonly encountered in current methods. The safety of the robot is guaranteed through the protective model described in Fig. 2(a). When the repulsion force from any obstacles is greater than the propulsion, then the robot must stop and switch to the fusion decision state for future planning directions:
$\exists_{B\text{ in }\Theta_{s}},(\vec{F}_{a}+\vec{F}_{aB})\cdot\vec{e}_{a}\leq 0.$ (14)
Subsequently, the cooperation and optimization are performed in Section III. Robots will search for a safe way towards the given direction. Then, it will circumnavigate the obstacle, and resumes to advance towards the target.
@@ IV-A Planning Process
In order to prevent path conflicts among robots, our model incorporates a collision avoidance state. When a robot is subjected to any repulsive interaction $\vec{F}_{ai}$ from neighbors greater than its own propulsion, it then enters the collision avoidance state:
@@
Figure 3: Finite state machine and its state transitions.
IV-B Parameter and Variable Determination
The proposed Nav-SCOPE is interpretable, as its parameters have explicit physical significance. We provide a way to assign parameters and variables without gaps for deployment.
IV-B1 Safety margin $r$
It is determined by robot’s max velocity $v_{m}$, max acceleration $a_{m}$, and sensor delay $t_{d}$:
@@ IV-B1 Safety margin $r$
where the first term refers to the radius of robot’s circumcircle, the second term represents the distance $s_{d}$ caused by sensor delay, and the third term denotes the braking distance $s_{b}$.
IV-B2 Reference distance $\sigma_{a}$
@@
IV-B3 Distance threshold $\mathbf{w}_{n}$
@@
IV-B4 Interaction force weight $w_{i}$
This weight in $\vec{f}_{ai}$ is critical for modulating the interaction amplitudes. It can be theoretically determined by AdaBoost [20] with error rates $\epsilon_{s}$ and $\epsilon_{n}$ of the corresponding probabilities $P_{a}$ and $P_{ni}$. Since Nav-SCOPE finely models the cooperation mechanism, $w_{i}$ can be obtained with robust performance using extremely small samples. This process is illustrated in Fig. 2(b). A sampling playground is randomly generated separate from the maps tested in simulations, and nine samples for each robot are used for an approximate estimate. Consequently, we get $\epsilon_{s}=0.245$ and $\epsilon_{n}=0.446$, and the optimal amplitude of $\vec{f}_{ai}$ in (11) should be:
@@ IV-B4 Interaction force weight $w_{i}$
The value of $w_{i}$ should be adjusted to approach the optimum value. We set $w_{i}=1$ in experimental applications and use $w_{i}=0.25,0.5,2.0,4.0$ in simulations for comparison.
IV-B5 Self observation $\mathbf{x_{0}}$
This vector records the estimated open-space area from $Q_{1}$ to $Q_{4}$. Incorporating environmental features described in (3), the elements of $\mathbf{x_{0}}$ correspond to the reconstructed polygonal area with each quadrant:
$x_{0,k}=\sum_{j=0}^{n_{k}}\frac{1}{2}d_{0,j}d_{0,j+1}\sin{\left(\phi_{0,j+1}-\phi_{0,j}\right)},k=1,\cdots,4$ (20)
where $n_{k}$ is the number of extrema within the $k$-th quadrant. The cases for $j=1$ to $n_{k}$ represent the extrema within the quadrant, while $j=0$ and $j=n_{i+1}$ are the cases at quadrant boundaries.
IV-B6 Neighbor observation $\mathbf{x}_{ni}$
@@
$x_{ni,k}=\begin{cases}\left(\max{d_{i,k}},\min{d_{i,k}}\right)/2&\text{local minima exist}\\ \max{d_{i,k}}&\text{other cases}\\ \end{cases},$ (21)
where $k=1,2$ for $Q_{1}$ and $Q_{2}$. Since local minima indicate potential obstacles, the minimum value delivers a message of potential obstructions. For the case of no local minima, the maxima value can effectively emphasize potential pathways.
V Simulation Results
@@
In this section, Nav-SCOPE is applied to robot swarms in cluttered and unknown environments on Gazebo using an i9-12900K CPU. We compare the performance of Nav-SCOPE with 3 baselines: a current reactive solution DH-Bug [21], a current SOTA method EGO-Swarm-v2 [4], and an ablation baseline $w_{i}=0$ referred to as Nav-I, concerning individual robots without neighboring scope to form the interaction fields. To evaluate performance, we use the following statistical metrics: (a) Robot arrival rate (AR). (b) Path redundancy $\gamma$, inspired by information theory, is defined as
@@
where $i$ is the robot index within a swarm, $l^{*}$ signifies the optimal path length, and $\bar{l}$ denotes the average path length. Here, we use the path from A* [22] with a resolution of 0.05m as the optimal path. (c) The duration of the task $t$. The maximum permitted time is set to 5 minutes.
Swarms of 15 robots are tested in very dense forests, about 0.35 obstacles$/m^{2}$. The physical characters of the robot are $r_{0}=$ 0.15m, $v_{m}=$0.5m/s, and $a_{m}=$ $\text{m}/\text{s}^{2}$. The sensor we used is a 5Hz LiDAR with an observation range of $R=$5m. The parameters are: $r=$ 0.30m, $l_{th}=$ 0.60m, $\alpha=$60, and $\sigma_{a}=$ 0.9m. We compared performance in three tasks:
@@
The statistical results for 20 random maps are detailed in Table I and Fig. 5. The optimal parameter $w_{i}$ also aligns with the sampled results in Section IV-B, as illustrated in Fig. 2(d).
Figure 5: Path redundancy across different tasks and algorithms.
Figure 6: Detailed cases of coordination. (a) Convergence. (b) Divergence.
@@
V-A Convergence effects
The arrow markings in Fig. 4(j)-(l) show some cases where paths are optimized through spontaneous synergies within local teams. A detailed case is depicted in Fig. 6(a). When a robot identifies a safe gap with fewer obstacles, it attracts nearby neighbor robots to follow its path. As a result, local teams are naturally formed by convergence effects along paths with relatively fewer obstacles. This type of effect is absent from the baseline methods. Specifically, the ablation study of Nav-I indicates that independent optimization has limited impacts on system performance, particularly in complicated tasks.
V-B Divergence effects
Note that the information from the cooperative perception may be negative when neighbors detect obstacles. Fig. 6(b) illustrates a case when robot B in front encountered an obstacle, the following robot A turned to the other side. Thus, robots can avoid unnecessary detours caused by perceptual limitations and optimize future paths with fused observation. EGO-Swarm-v2 also exhibited divergence effects as it tends to generate scattered paths to avoid collisions. But this function sometimes leads to unnecessary detours.
V-C Collision avoidance
This feature is achieved by the repulsive effect of the interaction field, where planning directions are considered to prevent future interferences between neighboring robots. Simulation results demonstrate safe and organized paths, especially when robots are enforced to avoid reciprocal collisions, as shown in Fig. 4(l). However, for EGO-Swarm-v2, the loss of DOF on the Z axis makes it challenging to find a collision-free path for mobile robots. Hence, planner timeouts and aggressive trajectories occasionally occurred, resulting in robot failures and longer task durations.
V-D Resource consumption
The statistical performances of computation and communication are summarized in Table II, where $t_{\text{plan}}$ denotes the planning time. Nav-SCOPE performs the same order of computation as minimal reactive methods and much lower communication compared to EGO-Swarm-v2. The computation mainly comes from the FFT on $m=1024$ points and the calculation of local extrema. Their time complexities are $O(m\log(m))$ and $O(m)$, respectively. The information flow with the swarm consumes a minimal bandwidth, which is further reduced by on-demand transmission. Consequently, Nav-SCOPE is an ultralight solution, suitable for deployment on robots with limited computational capacity or in scenarios with restricted communication.
@@ V-D Resource consumption
$t_{\text{plan}}$ Memory Bandwidth1 Package Size
@@
EGO-Swarm-v2 1.2ms 3.4MB 21.1Kbps 289.9B
@@
  • 1
    1 Each robot.
VI Real World Experiments
Besides simulations, we also performed real world experiments with a swarm of 5 robots. Each robot is equipped with an 8-core ARM v8.2 CPU and a 20Hz OS1 LiDAR with an effective range of 5m. Data transmission is performed by a 78.125 Kbps ad-hoc ultra wide band (UWB) network. The robots use AHD720P cameras and UWB ranging for mutual positioning, and employ CMP10A electronic compasses to determine relative orientation. The robot characters are: $r_{0}=$ 0.29m, $r_{e}=$ 0.43m, $v_{m}=$ 0.5m/s, $a_{m}=$ 4m/s2. Consequently, we set $r=$ 0.49m, $l_{th}=$ 0.83m, $\alpha=$ 72 and $\sigma_{a}=$ 1.09m. The LiDAR observation is represented as a polar coordinate array with approximately 360 points.

### v3

@@ IV-C3 Fusion optimization
where $dir=\{\text{left, right}\}$ for ground robots, $P_{dir}$ refers to the probability of exploration to the left, and $Z$ is a normalization factor. An illustrative example can be found in Fig. 4. The decision is based on the principle of Maximum A Posteriori (MAP) to address the subtle slope of the sigmoid function. The fusion process expands the field of view of individual robots, thus significantly reducing the uncertainty of environmental observations. Consequently, redundant paths are eliminated by convergence or divergence effects in terms of robot paths.
Figure 5: (a) Sampling process to determine the optimal $w_{i}$. (b) Magnitude curves of the interaction fields. (c) Finite state machine and its state transitions.
V Model Deployment
This section details the necessary steps for the deployment of Nav-SCOPE, including the planning process, the mutual positioning of swarm robots, and the determination of parameters.
Algorithm 1 Nav-SCOPE
1: Initialize neighbor indicators ${\mathcal{N}_{a}},\mathcal{R}_{a}\leftarrow\varnothing$
[…]
3: FFT points: $N_{f}\leftarrow 2^{\lceil\log_{2}M\rceil}$
[…]
7:for $R_{i}$ in swarm robots do
8:  if Received feature $\Sigma_{a}$ request then
9:   $\mathcal{R}_{a}$.add($R_{i}$)
10:  end if
11:  if $R_{i}$ is observed within the FOV then
12:   $\mathcal{N}_{a}$.add($R_{i}$)
13:   Obtain $\vec{r}_{ai}$ and calculate $\vec{f}_{ai}$ from (12)
14:  end if
15:end for/* Event-triggered information flow */
16:if $\mathcal{S}_{a}==\text{ii}$ then
17:  Request and receive features $\Sigma_{i}$ from $R_{i}\in\mathcal{N}_{a}$
18:  Calculate $P_{ni}$ from (14)
19:else
20:  $P_{ni}=1$
21:end if
22:if $\mathcal{S}_{a}==\text{ii}$ or $\mathcal{R}_{a}\neq\varnothing$ then
23:  $y_{a2}(m)\leftarrow\mathscr{F}^{-1}_{N_{f}}\left[\mathscr{F}_{N_{f}}[s_{a}(n)]\odot H_{2}\right]R_{M}(m)$
24:  $\Sigma_{a}\leftarrow\left\{(\phi_{a,j},d_{a,j})|\Delta y_{a2}(j)\Delta y_{a2}(j+1)<0\cap\Theta_{s}\right\}$
25:  Transmit $\Sigma_{a}$ to all the requested robots $R_{i}\in\mathcal{R}_{a}$
26:end if/* Force calculation for state transition */
27: Calculate $\vec{F}_{a},\vec{F}_{ai}$ from (8) and (13)
28:return  $\Theta_{s},\vec{F}_{a},\vec{F}_{ai}$
V-A Planning Process
The planning process based on Nav-SCOPE is a minimal reactive solution, which has proved to be versatile in a wide range of applications [7, 9, 3, 8]. A finite state machine (FSM) is illustrated in Fig. 5(c), where a robot $R_{a}$ will go through a series of states $\mathcal{S}_{a}\in\{\text{i, ii, iii, iv}\}$ planning towards the target. Algorithm 1 illustrates the detailed process of Nav-SCOPE that obtains safe directions and interaction forces for state transition, where $\mathscr{F}_{N_{f}}[\cdot]$ indicates the FFT transform with $N_{f}$ points, $\mathcal{N}_{a}$ is the group of observed neighbors, and $\mathcal{R}_{a}$ contains teammates that ask for environmental features. The information flow is formed only when needed.
During the planning process, the robot $R_{a}$ needs to update safe direction set (Lines 3-6) and observed neighbor list (Lines 7-15) continuously. When the target direction is within safe directions, that is, $\vec{e}_{a}\in\Theta_{s}$, then $R_{a}$ will advance straight towards the target ($\mathcal{S}_{a}=\text{i}$). Otherwise, the target is blocked by an obstacle $B$, and $R_{a}$ will request a cooperative perception ($\mathcal{S}_{a}=\text{ii}$, Lines 16-18) combined with its own observation (Lines 22-24) to find an optimal direction dirdir. A local group will be naturally formed with the neighboring robots $\mathcal{N}_{a}$, where the environmental features are sent (Lines 8-9 and Line 25 at neighbors) and gathered at $R_{a}$. Then, the interaction forces are calculated (Line 27), and the probabilistic fusion is achieved by (15). Afterwards, $R_{a}$ will search along the given dirdir towards a safe way closest to its target:
$\min_{m\in dir}\{\langle\theta_{m};\theta_{\vec{e}_{a}}\rangle\},\text{ s.t. }\theta_{m}\in\Theta_{s}.$ (16)
Subsequently, $R_{a}$ will enter the obstacle following state ($\mathcal{S}_{a}=\text{iii}$). In case there is a dead end, where $\Theta_{s}=\varnothing$, the robot will rotate in place towards dirdir while updating $\Theta_{s}$, before it meets $\Theta_{s}\neq\varnothing$ and finds a way out. It then keeps a safety margin from the obstacle borders, until it circumnavigates the obstacle, where $\vec{e}_{a}\in\Theta_{s}$ satisfies again. Finally, it resumes to advance towards the target. The reactive strategy of Nav-SCOPE diverges from PF methods [7, 8], as robot paths are generated directly within safe regions. This feature gets rid of complicated computing for repulsive fields, and avoids saddle points commonly encountered in current methods. Compared with Bug algorithms [9, 10], Nav-SCOPE introduces cooperative perception and fusion optimization, which achieve much better paths and coordination.
@@ V-A Planning Process
In order to prevent path conflicts among robots, our model incorporates a collision avoidance state ($\mathcal{S}_{a}=\text{iv}$). When a robot is subjected to any repulsive interaction $\vec{F}_{ai}$ from neighbors greater than its own propulsion, it then enters the collision avoidance state:
@@
V-B Mutual Positioning
In order to fuse cooperative perception within the same coordinate system, as well as calculate interaction forces, it is essential to determine relative poses and distances between neighboring robots. Instead of using external equipment, an engineering improvement is introduced to achieve mutual positioning in a distributed way. In this study, each robot is equipped with an ultra-wide band (UWB) beacon and four $90^{\circ}$ directional array antennas with three units, as shown in Fig. 4(d). The range and azimuth of neighboring robots are decided by the angle of arrival (AOA) of the UWB beacons and the time difference of arrival (TDOA) between antenna units [29]. The relative poses of robots can be determined from geomagnetism. In real-world experiments, these are achieved by Nooploop NIUB01 IOT models and CMP10A electronic compasses.
V-C Parameter and Variable Determination
The proposed Nav-SCOPE is interpretable, as its parameters are of explicit physical significance. We provide a way to assign parameters and variables without gaps for deployment.
V-C1 Safety margin $r$
It is determined by robot’s max velocity $v_{m}$, max acceleration $a_{m}$, and the sensor delay $t_{d}$:
@@ V-C1 Safety margin $r$
where the first term refers to the radius of robot’s circumcircle, the second term represents the distance $s_{d}$ caused by the sensor delay, and the third term denotes the braking distance $s_{b}$.
V-C2 Reference distance $\sigma_{a}$
@@
V-C3 Distance threshold $\mathbf{w}_{n}$
@@
V-C4 Interaction force weight $w_{i}$
It modulates the interaction amplitudes in $\vec{f}_{ai}$. The optimal value can be theoretically determined by AdaBoost [30] with error rates $\epsilon_{s}$ and $\epsilon_{n}$ of the corresponding probabilities $P_{a}$ and $P_{ni}$. If these parameters are not available, they can be estimated with robust performance using extremely small samples. A sampling playground is randomly generated separate from the maps in the simulation, and nine samples of the swarm are used for an approximate estimation. This process is illustrated in Fig. 5(a), and the estimated performances are shown in Fig. 7(b). Consequently, we get $\epsilon_{s}=0.245$ and $\epsilon_{n}=0.446$, and the optimal amplitude of $\vec{f}_{ai}$ in (15) should be:
@@ V-C4 Interaction force weight $w_{i}$
The value of $w_{i}$ should be adjusted to approach the optimal value. We set $w_{i}=1$ in experimental applications and use $w_{i}=0.25,0.5,2.0,4.0$ in simulations for comparison.
V-C5 Self observation $\mathbf{x_{0}}$
This vector records the estimated open space area from $Q_{1}$ to $Q_{4}$. Incorporating environmental features and linear interpolation described in Section IV-B, the elements of $\mathbf{x_{0}}$ correspond to the reconstructed area with each quadrant:
$x_{0,k}$ $=\frac{1}{2}\int^{\phi_{k,u}}_{\phi_{k,l}}\left(\frac{d_{k,u}-d_{k,l}}{\phi_{k,u}-\phi_{k,l}}\phi+\frac{d_{k,l}\phi_{k,u}-d_{k,u}\phi_{k,l}}{\phi_{k,u}-\phi_{k,l}}\right)^{2}d\phi$ (22)
$=\frac{\phi_{k,u}-\phi_{k,l}}{6}\left(d_{k,l}^{2}+d_{k,l}d_{k,u}+d_{k,u}^{2}\right),k=1,\cdots,4$
where $\phi_{k,u}$ and $\phi_{k,l}$ are the upper and lower boundaries of the $k$-th quadrant, and $d_{k,u},d_{k,l}$ are the corresponding distances.
V-C6 Neighbor observation $\mathbf{x}_{ni}$
@@
$x_{ni,k}=\begin{cases}\left(\max{d_{i,k}},\min{d_{i,k}}\right)/2&\exists\text{ local minima}\\ \max{d_{i,k}}&\text{other cases}\\ \end{cases},$ (23)
where $k=1,2$ for $Q_{1}$ and $Q_{2}$. Since local minima indicate potential obstacles, the minimum value delivers a message of potential obstruction. For the case of no local minima, the maxima value can effectively emphasize potential pathways.
VI Simulation Results
@@
In this section, Nav-SCOPE is applied to robot swarms in cluttered and unknown environments on Gazebo using an i9-12900K CPU. We compare the performance of Nav-SCOPE with 3 baselines: a current reactive solution DH-Bug [10], a current SOTA method EGO-Swarm-v2 [1], and an ablation baseline $w_{i}=0$ referred to as Nav-I, concerning individual robots without neighboring scope to form the interaction fields.
To evaluate performances, we use the following statistical metrics: (a) Robot arrival rate (AR), indicating the ratio of arrived robots to the total number of swarm robots. (b) Path redundancy $\gamma$, referring to the redundant path relative to the optimal path length. Inspired from the information theory [31], $\gamma$ is defined as
@@
where $i$ is the robot index within a swarm, $l^{*}$ signifies the optimal path length, and $\bar{l}$ denotes the average path length. Here, we use the path from A* [32] with a resolution of 0.05m as the optimal path. (c) Duration of the task $t$.
[…]
@@
The statistical results for 20 random maps are detailed in Table I and Fig. 7(a). The optimal parameter $w_{i}$ aligns with the sampled results in Section V-C, as illustrated in Fig. 7(b).
@@
Figure 7: (a) Path redundancy $\gamma$ across different tasks and algorithms. (b) Path redundancy $\gamma$ with different $w_{i}$.
VI-A Convergence effects
The arrow markings in Fig. 6(j)-(l) show some cases where paths are optimized through spontaneous synergies within local teams. A detailed case is depicted in Fig. 4(a). When a robot identifies a safe gap with fewer obstacles, it attracts nearby neighbor robots to follow its path. As a result, local teams are naturally formed by convergence effects along paths with relatively fewer obstacles. This type of effect is absent from the baseline methods. Specifically, the ablation study of Nav-I indicates that optimization with local perception alone has limited impacts on long-term performance, particularly in complicated tasks.
VI-B Divergence effects
Note that the information from the cooperative perception may be negative when neighbors detect obstacles. Fig. 4(b) illustrates a case when the robot $R_{b}$ in front encountered an obstacle, the following robot $R_{a}$ turned to the other side. Thus, robots can avoid unnecessary detours caused by perceptual limitations and optimize future paths with fused observation. EGO-Swarm-v2 also exhibited divergence effects as it tends to generate scattered paths to avoid collisions. But this function sometimes leads to unnecessary detours.
VI-C Collision avoidance
This feature is achieved by the repulsive effect of the interaction field, where planning directions are considered to prevent future interference between neighboring robots. Simulation results demonstrate safe and organized paths, especially when robots are enforced to avoid reciprocal collisions, as shown in Fig. 6(l). However, for EGO-Swarm-v2, the loss of DOF on the Z axis makes it challenging to find a collision-free path for mobile robots in dense situations. Hence, planner timeouts and aggressive trajectories occasionally occurred, resulting in robot failures and longer task durations.
@@ VI-C Collision avoidance
Nav-SCOPE 15.9 14.9 33.6 80.8
  • 1
    1 Syetem level.
VI-D Resource consumption
@@
VII Real World Experiments
Besides simulations, we also performed real-world experiments with a swarm of 5 robots. Each robot is equipped with an 8-core ARM v8.2 CPU and a 20Hz OS1 LiDAR with an effective range of 5m. Data transmission is performed by a 78.125 Kbps ad-hoc UWB network. The robots use Nooploop NIUB01 IOT models and CMP10A electronic compasses for mutual positioning. The robot characters are: $r_{0}=$ 0.29m, $r_{e}=$ 0.43m, $v_{m}=$ 0.5m/s, $a_{m}=$ 4m/s2. Consequently, we set $r=$ 0.49m, $l_{th}=$ 0.83m, $\alpha=$ 72 and $\sigma_{a}=$ 1.09m. The LiDAR observation is represented as a polar coordinate array with approximately 360 points.